L + x be a nœemal randome variable with mean 4 and variouse dot alpha Find the probability 8li * e_{s}




1) Between 4 and 6




2) Between 5 and 7.5




3) ・2 and 3.




4) greater them d * 75




5) Less than 4.025.



1
Expert's answer
2021-12-28T17:36:01-0500

mean:

"\mu=4"

standard deviation:

"\sigma=8"

"z=\frac{x-\mu}{\sigma}"


1)

"z_1=\frac{4-4}{8}=0"


"z_2=\frac{6-4}{8}=0.25"


"P(4<x<6)=P(z_1<z<z_2)=0.5987-0.5=0.987"


2)

"z_1=\frac{5-4}{8}=0.125"


"z_2=\frac{7.5-4}{8}=0.44"


"P(7.5<x<5)=P(z_1<z<z_2)=0.67-0.5478=0.1222"


3)

"z_1=\frac{2-4}{8}=-0.25"


"z_2=\frac{3-4}{8}=-0.125"


"P(2<x<3)=P(z_1<z<z_2)=0.4522-0.4013=0.0509"


4)

"z=\frac{7.5-4}{8}=0.44"


"P(x>7.5)=P(z>0.44)=1-0.67=0.33"


5)

"z=\frac{4.025-4}{8}=0.003"


"P(x<4.025)=P(z<0.003)=0.5"